# Quasirandom and quasisimple groups

### v2

@@ 4.2. Proof of Theorem B
From the data collected in the Appendix, we note that, among the finite simple groups, $J_{1}$ and $O^{\prime}N$ are the only ones whose quasirandom degree is at least $\frac{1}{3}$. Suppose that $G$ is of minimal cardinality among the finite groups satisfying $\bm{\varepsilon}(G)\geq\frac{1}{3}$ that are not isomorphic to either $J_{1}$ or $O^{\prime}N$. Of course, there exists a nontrivial proper normal subgroup $N$ of $G$. We have $\bm{\varepsilon}(G/N)\geq\frac{1}{3}$, and $G/N$ is isomorphic to either $J_{1}$ or $O^{\prime}N$. In the former case, from (2.1), we have
[…]
So $N=1$, a contradiction. Now suppose $G/N\cong O^{\prime}N$. From (2.1),
[…]
@@
We are ready to prove Theorem B. This proof has a number of cases, the most difficult, which we treat in the end, being the following. We have an abelian minimal normal subgroup $N$ of $G$ such that $G/N$ is quasisimple, and the action of $G/N$ on $N$ by conjugation provides a projective modular representation of the simple quotient of $G/N$, say $L$. The key point is that the (complex, linear) quasirandomness of $L$ controls also the degrees of the modular projective representations. This gives a lower bound on $|N|$, which we put against the upper bound given by (2.1). We start with the details.
@@ 2. $(b)$
If we are in case $(a)$, using (2.1) we obtain
[…]
Hence $N$ is trivial, a contradiction.
@@
(4.2) $|G|\>\leq\>\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}.$
Considering the embedding of $G$ into $\mathrm{Aut}(T^{n})$, and arguing as before, we obtain $\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))\leq n$. Since $|N|\geq 60^{n}$, and $60^{x}>x^{5}$ for every $x\geq 1$, this contradicts (4.2).
@@
Moreover, the action of $G/N$ on $N$ by conjugation provides a projective modular representation of $L$ (recall that $G/N$ is quasisimple). As already explained, the key idea is that the quasirandomness of $G/N$ gives a lower bound on $|N|$, which we put against the upper bound given by (2.1). We observe that $\mathrm{Alt}(5)$ and $\mathrm{Alt}(6)$ are $\frac{1}{4}$-quasirandom, but if one among them appears as $L$, then
[…]
@@ 3. $(iii)$
Let $L$ be as described in $(i)$. We remark that $|L|>\mathbf{D}_{0}(L)^{3}$ by Corollary D. Hence, by (2.1),
@@
Let $L$ be as described in $(ii)$. We first deal with $L\cong\mathrm{Sp}_{4}(q)$ with $q$ even, $q\geq 4$. Since the Schur multiplier is trivial, we have $N=M$. By (2.1), we find
@@
The argument to deal with the other cases is the same, using (2.1), Lemma 2.13 and Table 1. When $L\cong\mathrm{PSL}_{2}(q)$,
@@
Observe that $M_{11}$, $J_{1}$ and ${}^{2}F_{2}(4)^{\prime}$ admit no central perfect extensions, while the perfect central extensions of $J_{3}$ and $O^{\prime}N$ are not $\frac{1}{5}$-quasirandom. Therefore $N=M$ is the unique normal subgroup of $G$, and $N$ can be thought as the vector space of a nontrivial modular representation of $L$. For each sporadic group, we use [14] to obtain a lower bound on $|N|$, and we compare the result with (2.1). For $L\cong M_{11}$, we have
@@
Appendix: Quasirandom degree of finite simple groups
@@
In the following tables, we collect the values $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$ when $L$ is a finite simple group of Lie type or a sporadic simple group. For a group of Lie type $L_{d}(q)$, we fix $d$ and make $q$ go to infinity. The case where $L$ is of exceptional type is discussed in detail by Lübeck [17]. So let $L$ be a classical group of Lie type. In [27], Tiep and Zalesskii computed the minimal dimension of a projective complex irreducible representation of $L$, which we denote by $\mathbf{R}_{0}(L)$. See also [18]. In Table 2, we use these informations to record $\mathbf{D}_{0}(L)$ for the simple classical groups. We remark that $\mathbf{R}_{0}(L)<\mathbf{D}_{0}(L)$ holds in some cases, for example for some symplectic groups in odd characteristic.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d=2$, $q$ even $q-1$
[…]
@@
[…]
$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q\neq 3$ and $12(q^{2}-1)\nmid q^{2d}-$ 1 $\dfrac{q^{2d}-1}{q^{2}-1}$
$d\geq 3$, $q=3$ or $12(q^{2}-1)\mid q^{2d}-$ 1 $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
@@
Table 2. $\mathbf{D}_{0}(L)$ for simple classical groups.
In Tables 3, 4, and 5, we collect the asymptotic quasirandom degrees of the finite simple groups of Lie type, and the quasirandom degrees of the sporadic simple groups [7]. We remark that, after $(\mathrm{PSL}_{2}(q))_{q\geq 4}$, which is asymptotically $\frac{1}{3}$-quasirandom, the most quasirandom families of finite simple groups are $(\mathrm{Sp}_{4}(q))_{q\,even}$ and the Suzuki groups $(^{2}B_{2}(q^{2}))_{q\geq 8}$, whose quasirandom degrees are asymptotically $\frac{3}{10}$.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d\geq 2$ $d^{2}-1$ $d-1$ $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ $d^{2}-1$ $d-1$ $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSp}_{2d}(q)$ $d\geq 2$, $q$ even $2d^{2}+d$ $12d-$ 1 $\dfrac{2d-1}{2d^{2}+d}$
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q$ odd $2d^{2}+d$ $22d-$ 2 $\dfrac{2d-2}{2d^{2}+d}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{\pm}_{2d}(q)$ $d\geq 4$ $2d^{2}-d$ $32d-$ 3 $\dfrac{2d-3}{2d^{2}-d}$
Table 3. Asymptotic quasirandom degree of simple classical groups.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ odd $52$ $8$ $\dfrac{2}{13}$
[…]
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $10$ $3$ $\dfrac{3}{10}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $14$ $4$ $\dfrac{2}{7}$
Table 4. Asymptotic quasirandom degree of simple exceptional groups.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{11}$ 7 92079207\,9207 920 $10$ $\approx 0.2565$ $Fi_{24}$ 125 520 $125\,520\,570\cdot 10^{16}$ 8 67186718\,6718 671 $\approx 0.1635$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{12}$ 95 0409504095\,04095 040 $11$ $\approx 0.2093$ $HS$ 44 352 0004435200044\,352\,00044 352 000 $22$ $\approx 0.1756$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{22}$ 443 520443520443\,520443 520 $21$ $\approx 0.2342$ $McL$ 898 128 000898128000898\,128\,000898 128 000 $22$ $\approx 0.1500$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{23}$ 10 200 9601020096010\,200\,96010 200 960 $22$ $\approx 0.1916$ $He$ 4 030 387 20040303872004\,030\,387\,2004 030 387 200 $51$ $\approx 0.1778$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{24}$ 244 823 040244823040244\,823\,040244 823 040 $23$ $\approx 0.1624$ $Ru$ 145 926 $145\,926\,144\cdot 10^{3}$ $378$ $\approx 0.2309$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{1}$ 175 560175560175\,560175 560 $56$ $\approx 0.3334$ $Suz$ 448 345 $448\,345\,497\cdot 10^{3}$ $143$ $\approx 0.1850$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{2}$ 604 800604800604\,800604 800 $14$ $\approx 0.1924$ $O^{\prime}N$ 460 815 $460\,815\,505\cdot 10^{3}$ 10 9441094410\,94410 944 $\approx 0.3464$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{3}$ 50 232 9605023296050\,232\,96050 232 960 $85$ $\approx 0.2506$ $HN$ 273 030 $273\,030\,912\cdot 10^{6}$ $133$ $\approx 0.1472$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{4}$ 867 755 $867\,755\,710\cdot 10^{11}$ 1 33313331\,3331 333 $\approx 0.1568$ $Ly$ 517 651 $517\,651\,790\cdot 10^{8}$ 2 48024802\,4802 480 $\approx 0.2031$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{3}$ 495 766 $495\,766\,656\cdot 10^{3}$ $23$ $\approx 0.1165$ $Th$ 907 459 $907\,459\,438\cdot 10^{8}$ $248$ $\approx 0.1413$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{2}$ 423 054 $423\,054\,213\cdot 10^{5}$ $23$ $\approx 0.1000$ $B$ 415 478 $415\,478\,148\cdot 10^{24}$ 4 37143714\,3714 371 $\approx 0.1083$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{1}$ 415 777 $415\,777\,680\cdot 10^{10}$ $276$ $\approx 0.1311$ $M$ 808 017 $808\,017\,424\cdot 10^{44}$ 196 883196883196\,883196 883 $\approx 0.0983$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{22}$ 645 617 $645\,617\,516\cdot 10^{5}$ $78$ $\approx 0.1371$ ${}^{2}F_{4}(2)^{\prime}$ 17 971 2001797120017\,971\,20017 971 200 $104$ $\approx 0.2781$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{23}$ 408 947 $408\,947\,047\cdot 10^{10}$ $782$ $\approx 0.1555$
Table 5. Quasirandom degree of the sporadic simple groups.
Acknowledgments: We thank S. Eberhard for enlightening conversations on the paper, and P. Tiep for useful informations about his articles.

### v3

@@ 4.2 Proof of Theorem B
From the data collected in the Appendix, we note that, among the finite simple groups, $J_{1}$ and $O^{\prime}N$ are the only ones whose quasirandom degree is at least $\frac{1}{3}$. Suppose that $G$ is of minimal cardinality among the finite groups satisfying $\bm{\varepsilon}(G)\geq\frac{1}{3}$ that are not isomorphic to either $J_{1}$ or $O^{\prime}N$. Of course, there exists a nontrivial proper normal subgroup $N$ of $G$. We have $\bm{\varepsilon}(G/N)\geq\frac{1}{3}$, and $G/N$ is isomorphic to either $J_{1}$ or $O^{\prime}N$. In the former case, from (1), we have
$|N|\leq\frac{\mathbf{D}_{0}(J_{1})^{1/\bm{\varepsilon}(G)}}{|J_{1}|}\leq\frac{56^{3}}{175\,560}\approx 1.00031.$
So $N=1$, a contradiction. Now suppose $G/N\cong O^{\prime}N$. From (1),
$|N|\leq\frac{\mathbf{D}_{0}(O^{\prime}N)^{1/\bm{\varepsilon}(G)}}{|O^{\prime}N^{\prime}|}\leq\frac{10\,944^{3}}{460\,815\,505\,920}\approx 2.84446,$
@@
We are ready to prove Theorem B. This proof has a number of cases, the most difficult, which we treat in the end, being the following. We have an abelian minimal normal subgroup $N$ of $G$ such that $G/N$ is quasisimple, and the action of $G/N$ on $N$ by conjugation provides a projective modular representation of the simple quotient of $G/N$, say $L$. The key point is that the (complex, linear) quasirandomness of $L$ controls also the degrees of the modular projective representations. This gives a lower bound on $|N|$, which we put against the upper bound given by (1). We start with the details.
@@ 2. $(b)$
If we are in case $(a)$, using (1) we obtain
$|N|\leq\frac{\mathbf{D}_{0}((\mathbb{F}_{3})^{3}\rtimes\mathrm{SL}_{3}(3))^{5}}{|(\mathbb{F}_{3})^{3}\rtimes\mathrm{SL}_{3}(3)|}=\frac{12^{5}}{151\,632}\approx 1,64.$
Hence NN is trivial, a contradiction.
@@
$|G|\>\leq\>\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}.$ (3)
Considering the embedding of $G$ into $\mathrm{Aut}(T^{n})$, and arguing as before, we obtain $\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))\leq n$. Since $|N|\geq 60^{n}$, and $60^{x}>x^{5}$ for every $x\geq 1$, this contradicts (3).
@@
Moreover, the action of $G/N$ on $N$ by conjugation provides a projective modular representation of $L$ (recall that $G/N$ is quasisimple). As already explained, the key idea is that the quasirandomness of $G/N$ gives a lower bound on $|N|$, which we put against the upper bound given by (1). We observe that $\mathrm{Alt}(5)$ and $\mathrm{Alt}(6)$ are $\frac{1}{4}$-quasirandom, but if one among them appears as $L$, then
$|G|\leq\mathbf{D}_{0}(L)^{5}\leq 5^{5}=3\,125,$
@@ 3. $(iii)$
Let $L$ be as described in $(i)$. We remark that $|L|>\mathbf{D}_{0}(L)^{3}$ by Corollary D. Hence, by (1),
@@
Let $L$ be as described in $(ii)$. We first deal with $L\cong\mathrm{Sp}_{4}(q)$ with $q$ even, $q\geq 4$. Since the Schur multiplier is trivial, we have $N=M$. By (1), we find
@@
The argument to deal with the other cases is the same, using (1), Lemma 2.12 and Table 1. When $L\cong\mathrm{PSL}_{2}(q)$,
@@
Observe that $M_{11}$, $J_{1}$ and $^{2}F_{2}(4)^{\prime}$ admit no central perfect extensions, while the perfect central extensions of $J_{3}$ and $O^{\prime}N$ are not $\frac{1}{5}$-quasirandom. Therefore $N=M$ is the unique normal subgroup of $G$, and $N$ can be thought as the vector space of a nontrivial modular representation of $L$. For each sporadic group, we use [14] to obtain a lower bound on $|N|$, and we compare the result with (1). For $L\cong M_{11}$, we have
@@
Appendix
Appendix A Appendix: Quasirandom degree of finite simple groups
@@
In the following tables, we collect the values $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$ when $L$ is a finite simple group of Lie type or a sporadic simple group. For a group of Lie type $L_{d}(q)$, we fix $d$ and make $q$ go to infinity. The case where $L$ is of exceptional type is discussed in detail by Lübeck [17]. So let $L$ be a classical group of Lie type. In [27], Tiep and Zalesskii computed the minimal dimension of a projective complex irreducible representation of $L$, which we denote by $\mathbf{R}_{0}(L)$. See also [18]. In Table 2, we use this information to record $\mathbf{D}_{0}(L)$ for the simple classical groups. We remark that $\mathbf{R}_{0}(L)<\mathbf{D}_{0}(L)$ holds in some cases, for example for some symplectic groups in odd characteristic.
In Tables 3, 4, and 5, we collect the asymptotic quasirandom degrees of the finite simple groups of Lie type, and the quasirandom degrees of the sporadic simple groups [7]. We remark that, after $(\mathrm{PSL}_{2}(q))_{q\geq 4}$, which is asymptotically $\frac{1}{3}$-quasirandom, the most quasirandom families of finite simple groups are $(\mathrm{Sp}_{4}(q))_{q=2^{f}}$ and the Suzuki groups $(^{2}B_{2}(q^{2}))_{q\geq 8}$, whose quasirandom degrees are asymptotically $\frac{3}{10}$.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d=2$, $q$ even q1q-1
$d=2$, q=1(mod 4)q=1\,($\mathrm${mod}\,4) $\dfrac{q+1}{2}$
$d=2$, q=3(mod 4)q=3\,($\mathrm${mod}\,4) $\dfrac{q-1}{2}$
@@
$d\geq 2$, $q^{d}=1\,(\mathrm{mod}\,4)$ $\dfrac{q^{d}+1}{2}$
$d\geq 2$, $q^{d}=3\,(\mathrm{mod}\,4)$ $\dfrac{q^{d}-1}{2}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q\neq 3$ and $2(q^{2}-1)\nmid q^{2d}-1$ $\dfrac{q^{2d}-1}{q^{2}-1}$
$d\geq 3$, $q=3$ or $2(q^{2}-1)\mid q^{2d}-1$ $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
@@
Table 2: $\mathbf{D}_{0}(L)$ for simple classical groups.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d\geq 2$ $d^{2}-1$ d1d-1 $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ $d^{2}-1$ d1d-1 $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSp}_{2d}(q)$ $d\geq 2$, $q$ even $2d^{2}+d$ 2d12d-1 $\dfrac{2d-1}{2d^{2}+d}$
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q$ odd $2d^{2}+d$ 2d22d-2 $\dfrac{2d-2}{2d^{2}+d}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{\pm}_{2d}(q)$ $d\geq 4$ $2d^{2}-d$ 2d32d-3 $\dfrac{2d-3}{2d^{2}-d}$
Table 3: Asymptotic quasirandom degree of simple classical groups.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ odd $52$ 88 $\dfrac{2}{13}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $G_{2}(q)$ $q\neq 3\,(\mathrm{mod}\,6)$ $14$ 33 $\dfrac{3}{14}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $G_{2}(q)$ q=3(mod 6)q=3\,($\mathrm${mod}\,6) $14$ 44 $\dfrac{2}{7}$
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $10$ 33 $\dfrac{3}{10}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $14$ 44 $\dfrac{2}{7}$
Table 4: Asymptotic quasirandom degree of simple exceptional groups.
@@
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{11}$ 7 9207\,920 $10$ $\approx 0.2565$ $Fi_{24}$ $125\,520\,570\cdot 10^{16}$ 8 6718\,671 $\approx 0.1635$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{12}$ 95 04095\,040 $11$ $\approx 0.2093$ HSHS 44 352 00044\,352\,000 $22$ $\approx 0.1756$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{22}$ 443 520443\,520 2121 $\approx 0.2342$ McLMcL 898 128 000898\,128\,000 $22$ $\approx 0.1500$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{23}$ 10 200 96010\,200\,960 $22$ $\approx 0.1916$ HeHe 4 030 387 2004\,030\,387\,200 $51$ $\approx 0.1778$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{24}$ 244 823 040244\,823\,040 $23$ $\approx 0.1624$ RuRu $145\,926\,144\cdot 10^{3}$ $378$ $\approx 0.2309$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{1}$ 175 560175\,560 $56$ $\approx 0.3334$ SuzSuz $448\,345\,497\cdot 10^{3}$ $143$ $\approx 0.1850$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{2}$ 604 800604\,800 $14$ $\approx 0.1924$ $O^{\prime}N$ $460\,815\,505\cdot 10^{3}$ 10 94410\,944 $\approx 0.3464$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{3}$ 50 232 96050\,232\,960 $85$ $\approx 0.2506$ HNHN $273\,030\,912\cdot 10^{6}$ $133$ $\approx 0.1472$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{4}$ $867\,755\,710\cdot 10^{11}$ 1 3331\,333 $\approx 0.1568$ LyLy $517\,651\,790\cdot 10^{8}$ 2 4802\,480 $\approx 0.2031$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{3}$ $495\,766\,656\cdot 10^{3}$ $23$ $\approx 0.1165$ ThTh $907\,459\,438\cdot 10^{8}$ $248$ $\approx 0.1413$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{2}$ $423\,054\,213\cdot 10^{5}$ $23$ $\approx 0.1000$ $B$ $415\,478\,148\cdot 10^{24}$ 4 3714\,371 $\approx 0.1083$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{1}$ $415\,777\,680\cdot 10^{10}$ $276$ $\approx 0.1311$ $M$ $808\,017\,424\cdot 10^{44}$ 196 883196\,883 $\approx 0.0983$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{22}$ $645\,617\,516\cdot 10^{5}$ $78$ $\approx 0.1371$ $^{2}F_{4}(2)^{\prime}$ 17 971 20017\,971\,200 $104$ $\approx 0.2781$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{23}$ $408\,947\,047\cdot 10^{10}$ $782$ $\approx 0.1555$
Table 5: Quasirandom degree of the sporadic simple groups.
Acknowledgments
We thank an anonymous referee who found a mistake in Lemma 2.8. Moreover, we thank S. Eberhard and P.H. Tiep for useful conversations.